OpenStudy (anonymous):

Can someone please help me with this question? medal for anyone who helps :) If y varies inversely as the square of x, and y=1/8 when x=1, find y when x=5.

OpenStudy (luckycoins888):

um dats hard

OpenStudy (anonymous):

haha i know

OpenStudy (luckycoins888):

86 i tink

OpenStudy (anonymous):

...ill help

OpenStudy (anonymous):

how did you get 86 @luckycoins888 ? that's not one of the options for the question but i appreciate the help :)

OpenStudy (anonymous):

please help @crazychickengirl123 !!

OpenStudy (anonymous):

what are the options?

OpenStudy (anonymous):

that wold be d

OpenStudy (anonymous):

i meant d

OpenStudy (anonymous):

how did you get that though? i wanna know so i can do the next question myself :)

OpenStudy (anonymous):

internet

OpenStudy (anonymous):

use the internet

OpenStudy (anonymous):

i did...i couldn't find the exact answer.

OpenStudy (anonymous):

sorry but plz use the internet i have to go :(

OpenStudy (anonymous):

ok, thanks for the help!

OpenStudy (anonymous):

I'm sorry but @sleepyhead314 can help :)

OpenStudy (anonymous):

the statement "y varies directly as x" can be written as y = kx, where k is the variation constant we are given y=12 when x=2, so 12 = 2k k=6, so the variation constant is 6 and the equation is y = 6x

OpenStudy (sleepyhead314):

*sleepy sorry if I sound confusing* but it says inversely so y = k/x where k is a constant that we have to find

OpenStudy (sleepyhead314):

temporarily plug in y=1/8 and x= 1 for y and x into y = k/x and solve for k so 1/8 = k/1 solve for k

OpenStudy (sleepyhead314):

once you get k y = (k)/x plug in k then plug in x = 5 and solve for y

OpenStudy (sleepyhead314):

@crazychickengirl123 it says "inversely" not directly...

OpenStudy (anonymous):

shhhhhhh... im sleeping